Please help me with this problem :
Find the missing probability :
P(A) = 3/7, P(B) = 1/6. Find P( A and B)

Answer by stanbon(75887)

Please help me with this problem :
Find the missing probability :
To get an answer you have to assume that A and B are independent events:
------------------
P(A) = 3/7, P(B) = 1/6. Find P(A and B)
-----
P(A and B) = P(A)*P(B) = (3/7)(1/6) = 1/14
===============================================
Cheers,
Stan H.